# Channel Knowledge Map Construction via Physics-Inspired Diffusion Model Without Prior Observations

### v1

@@ III-B Denoising Diffusion Implicit Model
$\begin{split}q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t-1};\frac{\sqrt{\bar{\beta}_{t-1}-\sigma_{t}^{2}}}{\sqrt{\bar{\beta}_{t}}}\mathbf{x}_{t}\\ +\left(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\sqrt{\bar{\beta}_{t-1}-\sigma_{t}^{2}}}{\sqrt{\bar{\beta}_{t}}}\right)\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}).\end{split}$ (13)
@@
$\begin{split}q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\\ -\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}-\sigma_{t}^{2}})\boldsymbol{\epsilon}_{t},\sigma_{t}^{2}\mathbf{I}).\end{split}$ (14)
@@
$\begin{split}\mathbf{x}_{pre}=\sqrt{\bar{\alpha}_{pre}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}}{\sqrt{\bar{\alpha}_{t}}}\right)\\ +\sqrt{1-\bar{\alpha}_{pre}-\sigma_{t}^{2}}\boldsymbol{\epsilon}_{t}+\sigma_{t}^{2}\boldsymbol{\epsilon}.\end{split}$ (15)
@@
III-C Latent Diffusion Model
The diffusion model commonly employs a UNet architecture as the neural network for predicting either $\boldsymbol{\bar{\epsilon}}_{t}$ or $\mathbf{x}_{0}$. However, when the dimensionality of the input data is excessively high, the UNet inevitably becomes large in scale, leading to increased computational cost during both training and inference. The Latent Diffusion Model (LDM) first employs the encoder[45] first employs the encoder $\mathcal{E}$ of a Variational Autoencoder (VAE) to compress the high-dimensional input data $\mathbf{x}$ into a low-dimensional latent feature space $\mathbf{z}$. The UNet of the diffusion model is then trained in this latent space, which effectively reduces the network scale and computational cost. During the generation phase, the diffusion model produces a latent representation $\mathbf{z}_{0}$, which is subsequently decoded by the VAE decoder $\mathcal{D}$ to reconstruct the desired high-dimensional data $\mathbf{x}_{0}$.
In addition, the LDM introduces a novel conditional guidance mechanism that incorporates conditional prompts through a cross-attention module. This mechanism not only supports multimodal conditional inputs but also enables more precise conditional control, thereby providing a more flexible and effective approach to conditional guidance. The external conditional prompt $y$ is first encoded by the conditional encoder $\tau_{\theta}$ to obtain the $\tau_{\theta}(y)\in R^{d_{i}\times n}$. The internal feature tensor $\mathbf{z}_{t}$ is flattened to obtain the $\varphi(\mathbf{z}_{t})\in R^{d_{c}\times n}$. They are then multiplied by the matrices $W_{Q}\in R^{d_{n}\times d_{c}}$, $W_{K}\in R^{d_{n}\times d_{i}}$, $W_{V}\in R^{d_{n}\times d_{i}}$ to obtain $Q$, $K$, and $V$, respectively, as follows:
$Q=W_{Q}\cdot\varphi(\mathbf{z}_{t}),K=W_{K}\cdot\tau_{\theta}(y),V=W_{V}\cdot\tau_{\theta}(y).$ (17)
@@ III-C Latent Diffusion Model
IV The Physics-inspired Channel Knowledge Map Generation Diffusion Model
IV-A The Physical Meaning and Calculation Method of the Proposed Physical Regularization Terms
@@
IV-A1 Edge Loss Term
By observing the heatmap of the CGM, it can be found that its spatial distribution is highly consistent with an image captured under point-light illumination. Distinct edge structures appear in regions where obstacles cause signal occlusion. Therefore, as illustrated in Fig. 2, edge extraction can be applied to the CGM, and the resulting edge map depicts partial propagation paths of wireless signals after interacting with obstacles such as buildings. The residual between the edge map of the real CGM and that of the generated CGM is defined as the edge loss. This loss term ensures that the propagation paths from the base station to various spatial locations in the generated CGM are consistent with the surrounding environment and conform to the physical constraint that large-scale wireless signal attenuation is primarily dominated by the LoS component.
@@ IV-A1 Edge Loss Term
There exist various methods for edge detection. Considering computational efficiency, we adopt the Laplace operator to perform edge extraction on the CGM. The discrete form of the Laplace operator used for edge extraction is expressed as follows:
$\begin{split}\nabla^{2}\mathbf{x}(i,j)=\frac{\mathbf{x}(i+1,j)+\mathbf{x}(i-1,j)+\mathbf{x}(i,j+1)}{h^{2}}\\ +\frac{\mathbf{x}(i,j-1)-4\mathbf{x}(i,j)}{h^{2}},\end{split}$ (20)
where h represents the grid interval. After applying the Laplace operator to the CGM, a binary edge map can be obtained by thresholding the result. However, in deep learning, the binarization operation blocks gradient backpropagation, preventing the model from being trained end-to-end. To address this issue, we employ an approximate gradient propagation technique known as the Straight-Through Estimator (STE) to compute the edge loss term. When the input is denoted as $\mathbf{x}$, the STE function is defined as follows:
@@
where $\mathbf{x}_{0,n}$ denotes the real CGM samples, $\hat{\mathbf{x}}_{0,n}$ represents the CGM samples generated during training, and $N$ indicates the total number of samples.
IV-A2 Regional Propagation Loss Term
If there are no obstacles within a region, the path loss increases logarithmically with propagation distance. However, in the presence of obstacles, the spatial distribution of the CGM becomes nonuniform and strongly correlated with the locations of these obstacles. By examining the heatmap of the CGM, it can be observed that the map is segmented into multiple color blocks due to the influence of buildings and other obstructions. Areas with no obstacles between the base station and the receiver appear as lighter color regions, indicating smaller path losses, whereas regions with multiple obstacles appear darker, representing larger path losses.
@@ IV-A2 Regional Propagation Loss Term
In this work, a neural network is employed to directly map the CGM to its corresponding regional propagation map. As described earlier, the process of regional propagation division is analogous to image segmentation in computer vision. Therefore, we adopt the UNet architecture—renowned for its strong segmentation capability—to perform this task, referred to here as RegionUNet. During training, the target regional maps in the dataset are pre-obtained using the aforementioned multi-threshold division method. After training, the network parameters of RegionUNet are fixed, and the network is then used during the diffusion model training stage to compute the regional propagation maps of both the real and generated CGMs, as illustrated in Fig. 4.
@@
IV-A3 Multi-scale Feature Loss Term
@@
The lower layers capture fine-scale variations, such as the rapid signal attenuation at building edges or narrow streets caused by local shadowing and diffraction effects. The middle layers reflect medium-scale variations, representing path loss changes across different urban blocks and propagation behaviors within urban canyons. The upper layers characterize large-scale trends, describing the overall propagation tendency of the signal, i.e., large-scale fading in free-space environments. Therefore, by applying the Laplacian pyramid to both the real and generated CGMs, we extract their multi-scale feature maps, and the residual between these feature maps is incorporated into the training process as a physics-inspired regularization term. This design ensures that the generated CGM conforms to the multi-scale physical effects inherent in wireless signal propagation. The procedure for multi-scale feature extraction using the Laplacian pyramid is illustrated in Fig. 5.
@@
The input to the first layer of the Laplacian pyramid is the CGM. At layer $l$, the input is denoted by $\mathbf{x}_{l}$, the output $\mathbf{x}_{l+1}$ is obtained by applying Gaussian smoothing to $\mathbf{x}_{l}$ followed by $2\times$ downsampling. The feature map at scale $l$ is then computed as the difference between $\mathbf{x}_{l}$ and the $2\times$ upsampled version of $\mathbf{x}_{l+1}$. Therefore, the multi-scale feature loss term is computed by performing Laplacian pyramid decomposition, as shown in Fig. 5, on both the real and generated CGMs. The residuals between the corresponding feature maps at each scale are then summed to obtain the overall loss. The calculation formula is expressed as follows:
@@
IV-B The Derivation of Model Loss Function
Discriminative models, such as autoencoders, UNet, and LSTM, establish a direct mapping from input to output; therefore, their loss functions are typically defined as the Mean Squared Error (MSE) between the input and the output. When constructing a physics-inspired model based on a discriminative framework, the physics-inspired terms can be directly appended to the original MSE loss function, since both components jointly optimize the mapping relationship from input to output. However, as described in Section III, Subsection A, the neural network $p_{\theta}(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t})$ in diffusion models is designed to approximate the true posterior distribution $q(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t},\mathbf{x}_{0})$. Although, under the $\text{pred}-\mathbf{x}_{0}$ mode, the loss function can be simplified to the MSE between the network output and the input sample in form, the neural network in this mode still aims to fit the mean of the posterior distribution. Therefore, its output cannot be directly used as the generated sample. Since the optimization objective of the loss function in the vanilla diffusion model does not align with that required by our designed physics-inspired regularization terms, the loss function formulation of the physics-guided diffusion model needs to be rederived.
@@ IV-B The Derivation of Model Loss Function
Then, we define a virtual observation $\tilde{r}$ to introduce uncertainty into the model’s physical residual $r(\hat{\mathbf{x}}_{0})$[46]. The distribution of $\tilde{r}$ is given as follows:
@@
$\arg\max_{\theta}E_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0})}\left[\log p_{\theta}(\mathbf{x}_{0})\right]$ (28)
$+E_{\mathbf{x}_{1:T}\sim q(\mathbf{x}_{1:T})}\left[\log q_{R}(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{1:T}))\right]$
$=$ $\arg\min_{\theta}E_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0})}\left[-\log p_{\theta}(\mathbf{x}_{0})\right]$
$+E_{\mathbf{x}_{1:T}\sim q(\mathbf{x}_{1:T})}\left[-\log q_{R}(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{1:T}))\right].$
Based on the construction of the variational lower bound in[43] for deriving the DDPM loss function, the numerical formulation of the loss function for the proposed physics-inspired diffusion model can be derived as follows:
$E_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]$ (29)
$+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$
$\leq$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]$
$+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(x_{1:T}\right)\right)\right]$
@@
$=$ $E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}$
$+\frac{1}{2\sigma^{2}}\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}].$
The detailed derivation of Eq.(29) is provided in the Appendix, where $\tilde{A}_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}$, $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$. Similar to[43], we observe that ignoring the preceding weights leads to better generative performance. Therefore, the final loss function is formulated as follows:
$L_{\text{phy}}(\theta)=E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}$ (30)
$+\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}].$
IV-C The Physics-inspired Channel Knowledge Map Generation Diffusion Model Architecture and Algorithm Flow
The overall training framework of the proposed physics-inspired generation diffusion model is illustrated in Fig. (6). As shown in Fig. (6a), the first stage trains the RegionUNet network, which is employed to compute the regional propagation loss term. The second stage involves training a VAE that compresses the input data into a low-dimensional latent tensor, and the output of the VAE is defined as $\bar{\mathbf{x}}_{0}$, as depicted in Fig. (6b). After the completion of the first two stages, the network parameters of both the RegionUNet and the VAE are frozen and excluded from the training process in the third stage. The training framework of the diffusion model in the third stage is illustrated in Fig. (6c). The conditional prompt $\mathbf{c}$ includes the building layout map, the base station location map, and the vehicle location map. For the conditional encoder, we employ the Swin-T (Shifted Window Transformer) network, which has demonstrated strong performance in image processing tasks. The conditional prompts are encoded by the Swin-T and injected into the main UNet backbone through a cross-attention mechanism. The UNet backbone and the Swin-T encoder are jointly trained. During training, in addition to the original loss term $L_{z_{0}}$ defined in the LDM, three physics-inspired regularization terms, $L_{\text{edge}}$, $L_{\text{region}}$, $L_{\text{mul-fea}}$, are also incorporated into the total loss function. Therefore, unlike other diffusion models, each computation of the total loss function $L_{\text{total}}$ requires performing a reverse denoising process to generate the CGM. However, employing the reverse denoising sampling procedure of the DDPM to obtain the generated CGM would result in an unacceptably long training time. To address this issue, we adopt the DDIM during training to generate CGM $\hat{\mathbf{x}}_{0}$. The algorithmic procedure for DDIM sampling under the $\text{pred}-\mathbf{x}_{0}$ mode is presented in Algorithm 1.

### v2

@@ 3.2 Denoising Diffusion Implicit Model
$\begin{split}q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t-1};\frac{\sqrt{\bar{\beta}_{t-1}-\sigma_{t}^{2}}}{\sqrt{\bar{\beta}_{t}}}\mathbf{x}_{t}+\left(\sqrt{\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t}}\sqrt{\bar{\beta}_{t-1}-\sigma_{t}^{2}}}{\sqrt{\bar{\beta}_{t}}}\right)\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}).\end{split}$ (13)
@@
$\begin{split}q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}-\sigma_{t}^{2}})\boldsymbol{\epsilon}_{t},\sigma_{t}^{2}\mathbf{I}).\end{split}$ (14)
@@
$\begin{split}\mathbf{x}_{pre}=\sqrt{\bar{\alpha}_{pre}}\left(\frac{\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{pre}-\sigma_{t}^{2}}\boldsymbol{\epsilon}_{t}+\sigma_{t}^{2}\boldsymbol{\epsilon}.\end{split}$ (15)
@@
3.3 Latent Diffusion Model
The diffusion model commonly employs a UNet architecture as the neural network for predicting either $\boldsymbol{\bar{\epsilon}}_{t}$ or $\mathbf{x}_{0}$. However, when the dimensionality of the input data is excessively high, UNet inevitably becomes large in scale, leading to increased computational cost during both training and inference. The Latent Diffusion Model (LDM)[Rombach2021HighResolutionIS] first employs the encoder $\mathcal{E}$ of a Variational Autoencoder (VAE) to compress the high-dimensional input data $\mathbf{x}$ into a low-dimensional latent feature space $\mathbf{z}$. The UNet of the diffusion model is then trained in this latent space, which effectively reduces the network scale and computational cost. During the generation phase, the diffusion model produces a latent representation $\mathbf{z}_{0}$, which is subsequently decoded by the VAE decoder $\mathcal{D}$ to reconstruct the desired high-dimensional data $\mathbf{x}_{0}$.
In addition, the LDM introduces a novel conditional guidance mechanism that incorporates conditional prompts through a cross-attention module. This mechanism not only supports multimodal conditional inputs but also enables more precise conditional control, thereby providing a more flexible and effective approach to conditional guidance. The external conditional prompt $y$ is first encoded by the conditional encoder $\tau_{\theta}$ to obtain $\tau_{\theta}(y)\in\mathbb{R}^{n\times d_{i}}$. The internal feature tensor $\mathbf{z}_{t}$ is flattened to obtain the $\varphi(\mathbf{z}_{t})\in\mathbb{R}^{n\times d_{c}}$. They are then multiplied by the matrices $W_{Q}\in\mathbb{R}^{d_{c}\times d_{n}}$, $W_{K}\in\mathbb{R}^{d_{i}\times d_{n}}$, $W_{V}\in\mathbb{R}^{d_{i}\times d_{n}}$ to obtain $Q$, $K$, and $V$, respectively, as follows:
$Q=\varphi(\mathbf{z}_{t})\cdot W_{Q},K=\tau_{\theta}(y)\cdot W_{K},V=\tau_{\theta}(y)\cdot W_{V}.$ (17)
@@ 3.3 Latent Diffusion Model
4 Physics-inspired Channel Knowledge Map Generation Diffusion Model
4.1 The Physical Meaning and Calculation Method of the Proposed Physical Regularization Terms
@@
4.1.1 Edge Loss Term
By observing the heatmap of the CGM, it can be found that its spatial distribution is highly consistent with an image captured under point-light illumination. Distinct edge structures appear in regions where obstacles cause signal occlusion. Therefore, as illustrated in Fig. 2, edge extraction can be applied to the CGM, and the resulting edge map depicts partial propagation paths of wireless signals after interacting with obstacles such as buildings.
@@ 4.1.1 Edge Loss Term
The residual between the edge map of the ground truth CGM and that of the generated CGM is defined as the edge loss. This loss term ensures that the propagation paths from the BS to various spatial locations in the generated CGM are consistent with the surrounding environment and conform to the physical constraint that large-scale wireless signal attenuation is primarily dominated by the LoS component.There exist various methods for edge detection. Considering computational efficiency, we adopt the Laplace operator to perform edge extraction on the CGM. The discrete form of the Laplace operator used for edge extraction is expressed as follows:
$\begin{split}\nabla^{2}\mathbf{x}(i,j)=\frac{\mathbf{x}(i+1,j)+\mathbf{x}(i-1,j)+\mathbf{x}(i,j+1)}{h^{2}}+\frac{\mathbf{x}(i,j-1)-4\mathbf{x}(i,j)}{h^{2}},\end{split}$ (20)
where $h$ represents the grid interval. After applying the Laplace operator to the CGM, a binary edge map can be obtained by thresholding the result. However, in deep learning, the binarization operation blocks gradient backpropagation, preventing the model from being trained end-to-end. To address this issue, we employ an approximate gradient propagation technique known as the Straight-Through Estimator (STE)[bengio2013estimating] to compute the edge loss term. When the input is denoted as $\mathbf{x}$, the STE function is defined as follows:
@@
where $\mathbf{x}_{0,n}$ denotes the ground truth CGM samples, $\hat{\mathbf{x}}_{0,n}$ represents the CGM samples generated during training, and $N$ indicates the total number of samples.
4.1.2 Regional Propagation Loss Term
If there are no obstacles within a region, the path loss increases logarithmically with propagation distance. However, in the presence of obstacles, the spatial distribution of the CGM becomes nonuniform and strongly correlated with the locations of these obstacles. By examining the heatmap of the CGM, it can be observed that the map is segmented into multiple color blocks due to the influence of buildings and other obstructions. Areas with no obstacles between the BS and the receiver appear as lighter color regions, indicating smaller path losses, whereas regions with multiple obstacles appear darker, representing larger path losses.
@@ 4.1.2 Regional Propagation Loss Term
In this work, a neural network is employed to directly map the CGM to its corresponding regional propagation map. As described earlier, the process of regional propagation division is analogous to image segmentation in computer vision. Therefore, we adopt the UNet architecture—renowned for its strong segmentation capability—to perform this task, referred to here as RegionUNet. During training, the target regional maps in the dataset are pre-obtained using the aforementioned multi-threshold division method. After training, the network parameters of RegionUNet are fixed, and the network is then used during the diffusion model training stage to compute the regional propagation maps of both the ground truth and generated CGMs, as illustrated in Fig. 4.
@@
4.1.3 Multi-scale Feature Loss Term
@@
The lower layers capture fine-scale variations, such as the rapid signal attenuation at building edges or narrow streets caused by local shadowing and diffraction effects. The middle layers reflect medium-scale variations, representing path loss changes across different urban blocks and propagation behaviors within urban canyons. The upper layers characterize large-scale trends, describing the overall propagation tendency of the signal, i.e., large-scale fading in free-space environments. Therefore, by applying the Laplacian pyramid to both the ground truth and generated CGMs, we extract their multi-scale feature maps, and the residual between these feature maps is incorporated into the training process as a physics-inspired regularization term. This design ensures that the generated CGM conforms to the multi-scale physical effects inherent in wireless signal propagation. The procedure for multi-scale feature extraction using the Laplacian pyramid is illustrated in Fig. 5.
@@
The input to the first layer of the Laplacian pyramid is the CGM. At layer $l$, the input is denoted by $\mathbf{x}_{l}$, the output $\mathbf{x}_{l+1}$ is obtained by applying Gaussian smoothing to $\mathbf{x}_{l}$ followed by $2\times$ downsampling. The feature map at scale $l$ is then computed as the difference between $\mathbf{x}_{l}$ and the $2\times$ upsampled version of $\mathbf{x}_{l+1}$. Therefore, the multi-scale feature loss term is computed by performing Laplacian pyramid decomposition, as shown in Fig. 5, on both the ground truth and generated CGMs. The residuals between the corresponding feature maps at each scale are then summed to obtain the overall loss. The calculation formula is expressed as follows:
@@
4.2 The Derivation of Model Loss Function
Discriminative models, such as autoencoders, UNet, and LSTM, establish a direct mapping from input to output; therefore, their loss functions are typically defined as the Mean Squared Error (MSE) between the input and the output. When constructing a physics-inspired model based on a discriminative framework, the physics-inspired terms can be directly appended to the original MSE loss function, since both components jointly optimize the mapping relationship from input to output. However, as described in Subsection III A, the neural network $p_{\theta}(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t})$ in diffusion models is designed to approximate the true posterior distribution $q(\mathbf{x}_{t-1}\!\mid\!\mathbf{x}_{t},\mathbf{x}_{0})$. Although, under the $\text{pred}-\mathbf{x}_{0}$ mode, the loss function can be simplified to the MSE between the network output and the input sample in form, the neural network in this mode still aims to fit the mean of the posterior distribution. Therefore, its output cannot be directly used as the generated sample. Since the optimization objective of the loss function in the vanilla diffusion model does not align with that required by our designed physics-inspired regularization terms, the loss function formulation of the physics-guided diffusion model needs to be rederived.
@@ 4.2 The Derivation of Model Loss Function
Then, we define a virtual observation $\tilde{r}$ to introduce uncertainty into the model’s physical residual $r(\hat{\mathbf{x}}_{0})$[RIXNER2021110218]. The distribution of $\tilde{r}$ is given as follows:
@@
$\arg\max_{\theta}E_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0})}\left[\log p_{\theta}(\mathbf{x}_{0})\right]+E_{\mathbf{x}_{1:T}\sim q(\mathbf{x}_{1:T})}\left[\log q_{R}(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{1:T}))\right]$ (28)
$=$ $\arg\min_{\theta}E_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0})}\left[-\log p_{\theta}(\mathbf{x}_{0})\right]+E_{\mathbf{x}_{1:T}\sim q(\mathbf{x}_{1:T})}\left[-\log q_{R}(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}(\mathbf{x}_{1:T}))\right].$
Based on the construction of the variational lower bound in[ho2020denoising] for deriving the DDPM loss function, the numerical formulation of the loss function for the proposed physics-inspired diffusion model can be derived as follows:
$E_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left[-\log p_{\theta}\left(\mathbf{x}_{0}\right)\right]+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{1:T}\right)\right)\right]$ (29)
$\leq$ $E_{q\left(\mathbf{x}_{0:T}\right)}\left[\log\frac{q\left(\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\right)}{p_{\theta}\left(\mathbf{x}_{0:T}\right)}\right]+E_{\mathbf{x}_{1:T}\sim q\left(\mathbf{x}_{1:T}\right)}\left[-\log q_{R}\left(\tilde{r}=0\mid\hat{\mathbf{x}}_{0}\left(x_{1:T}\right)\right)\right]$
@@
$=$ $E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\frac{\tilde{A}_{t}^{2}}{2\tilde{\beta}_{t}^{2}}\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}+\frac{1}{2\sigma^{2}}\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}].$
The detailed derivation of Eq.(29) is provided in the Appendix, where $\tilde{A}_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}$, $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$. Similar to[ho2020denoising], we observe that ignoring the preceding weights leads to better generative performance. Therefore, the final loss function is formulated as follows:
$L_{\text{phy}}(\theta)=E_{t\sim[1,T],\mathbf{x}_{0:T}\sim q\left(\mathbf{x}_{0:T}\right)}[\left\|\mathbf{x}_{0}-\mathbf{x}_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}+\left\|r\left(\hat{\mathbf{x}}_{0}\left(\mathbf{x}_{t},t\right)\right)\right\|^{2}].$ (30)
@@
4.3 The Physics-inspired Channel Knowledge Map Generation Diffusion Model Architecture and Algorithm Flow
The overall training framework of the proposed physics-inspired generation diffusion model is illustrated in Fig. (6). As shown in Fig. (6a), the first stage trains the RegionUNet network, which is employed to compute the regional propagation loss term. The second stage involves training a VAE that compresses the input data into a low-dimensional latent tensor, and the output of the VAE is defined as $\bar{\mathbf{x}}_{0}$, as depicted in Fig. (6b). After the completion of the first two stages, the network parameters of both the RegionUNet and the VAE are frozen and excluded from the training process in the third stage. The training framework of the diffusion model in the third stage is illustrated in Fig. (6c). The conditional prompt $\mathbf{c}$ includes the building layout map, the BS location map, and the vehicle location map. For the conditional encoder, we employ the Swin-T (Shifted Window Transformer) network, which has demonstrated strong performance in image processing tasks. The conditional prompts are encoded by the Swin-T and injected into the main UNet backbone through a cross-attention mechanism. The UNet backbone and the Swin-T encoder are jointly trained. During training, in addition to the original loss term $L_{z_{0}}$ defined in the LDM, three physics-inspired regularization terms, $L_{\text{edge}}$, $L_{\text{region}}$, $L_{\text{mul-fea}}$, are also incorporated into the total loss function. Therefore, unlike other diffusion models, each computation of the total loss function $L_{\text{total}}$ requires performing a reverse denoising process to generate the CGM. However, employing the reverse denoising sampling procedure of the DDPM to obtain the generated CGM would result in an unacceptably long training time. To address this issue, we adopt the DDIM during training to generate CGM $\hat{\mathbf{x}}_{0}$. The algorithmic procedure for DDIM sampling under the $\text{pred}-\mathbf{x}_{0}$ mode is presented in Algorithm 1.